OpenStudy (anonymous):

Question: 48 of 63: James purchased 2 T-shirts and a pair of jeans online and paid $80 for his purchase. The next week, the same site was offering a 50% discount on all its clothes. James bought 3 T-shirts and two pairs of jeans for $70. What were the original prices of T-shirts and jeans? (Assume that all T-shirts had the same original price, as did all pairs of jeans.)

OpenStudy (anonymous):

Select one of the options below as your answer: A. T-shirt, $25; jeans, $30 B. T-shirt, $20; jeans, $30 C. T-shirt, $20; jeans, $40 D. T-shirt, $25; jeans, $35

OpenStudy (anonymous):

Let price of T-shirt be T, price of jeans be J. 2T + J = 80 ---- (1) After 50% discount, 3(0.5T) + 2(0.5J) = 70 ----- (2) J = 80 - 2T (from 1) sub J into 2 1.5T + 80 - 2T = 70 10 = 0.5T T = 20 thus J = 80 - 2(20) = 40